Calculus I, Fall 2012 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 10

1. Section 6.1: Problems 45, 47, 59, 65.

2. Section 6.2: Problems 28(c), 32, 45, 46.

3. Section 6.3: Problems 41, 50, 51, 57, 65, 77, 87, 89, 91.

4. Section 6.3: Do as many as time permits in problems 31-40, 85-90.

5. Section 6.4: Problems: 9, 19, 25, 41, 47, 57, 71.

6. (s6.4-extra1) Evaluate

lim_{n→∞} ∑_{k= \frac{n}{2}}^{n} \frac{1}{k}

Hint: try to express it in terms of \frac{k}{n}.